# Evaluating the Ebb and Flow: An In-depth Analysis of Question-Answering Trends across Diverse Platforms

### 08 Dec 2023 v3

@@ Abstract
Community Question Answering (CQA) platforms steadily gain popularity as they provide users with fast responses to their queries. The swiftness of these responses is contingent on a mixture of query-specific and user-related elements. This paper scrutinizes these contributing factors within the context of six highly popular CQA platforms, identified through their standout “answering speed”. Our investigation reveals a correlation between the time taken to yield the first response to a question and several variables: the metadata, the formulation of the questions, and the level of interaction among users. Additionally, by employing conventional machine learning models to analyze these metadata and patterns of user interaction, we endeavor to predict which queries will receive their initial responses promptly.
Keywords:
community question answering response time asker-answerer graph classification model
@@ 1 Introduction
Community question-answering platforms are gradually evolving over the past decade. Such portals allow users to post their queries and answer the questions. These CQA platforms help experts share their knowledge and new users solve their queries. Initially, these community question-answering platforms focused on providing the most relevant answers to the users’ queries [14][2]. As the CQA platforms keep growing, they not only focus on providing the relevant answers but also on quickly delivering such answers [12][11]. The growth and maturity of such platforms is contingent on a number of variables, including the “quality of questions and answers”, the “response time of queries”, “user involvement”, and “user activity” [8]. The users with expertise in specific domains significantly help the platform’s growth. Based on the users’ activeness on the platform and the importance of their posts, they have been rewarded reputation scores, badges, and additional privileges. In most such CQA platforms driven by a reputation system, the platform’s popularity depends on the response time [3] of the question (i.e., time to receive the first answer). So, the users who want to post a query may also want to know the time by which they can expect a response to their question. StackExchange [^1] https://stackexchange.com/ is one of the CQA platforms where the users can ask questions and receive answers from other users. An analysis of such a platform’s dynamics helps the community maintainers adapt and design the framework to address various CQA-related problems [1] [7] [13] [9] [6] [5] [4]. In this paper, we examine the dynamics of different CQA platforms relating the usual response time with other factors of the portals. We conduct an empirical study on six CQA platforms – Mathematics, Software Engineering, English, Game Development, Chemistry, and Money. To analyze the dynamics of the diverse community question-answering platforms, we look into various factors mentioned here – (1) Characterize the questions in terms of textual information, linguistic characteristics, question tags, votes received by the questions etc. We obtain correlations of these features with the time required to respond to a question. (2) Characterizing the users of the platform in terms of their reputation & activity in the platforms. Further, we build Asker-Answerer graph (AAG) where the nodes are the users and a link is created if the question from a user is answered by another user. Subsequently, we conduct in depth analysis of the characteristics of the common users present across all the platforms. (3)For all the platforms, we use metadata of the questions, question structure and user interaction as features to build a standard machine learning model for classifying the newly posted questions which are having fast answer. The majority of preceding studies [10] scrutinize the correlation between diverse question-related factors and response time. The novelty of our current research resides in our endeavor to comprehend how various inherent network-related features (reflected in the asker-answerer graph) coincide with question-related features in relation to response time. In this study, we make the following observations – (1) Short questions get faster responses. Chemistry and Software Engineering questions are harder. More tags mean slower replies. Math, Software Engineering, and English answers come in an hour, others in a day. In Software Engineering and English, many questions get $\geq 3$ answers.(2) Most users ( 60%-70%) on all platforms stay inactive. Domains like Money, Game Development, and Chemistry have more active users than Mathematics, Software Engineering, and English. (3) XGBoost and MLP perform best at classifying new questions on most platforms. The second best is usually the random forest model.

### 15 Mar 2024 v5

@@ Abstract
Community Question Answering (CQA) platforms steadily gain popularity as they provide users with fast responses to their queries. The swiftness of these responses is contingent on a mixture of query-specific and user-related elements. This paper scrutinizes these contributing factors within the context of six highly popular CQA platforms, identified through their standout “answering speed". Our investigation reveals a correlation between the time taken to yield the first response to a question and several variables: the metadata, the formulation of the questions, and the level of interaction among users. Additionally, by employing conventional machine learning models to analyze these metadata and patterns of user interaction, we endeavor to predict which queries will receive their initial responses promptly.
@@ 1 Introduction
Community question-answering platforms are gradually evolving over the past decade. Such portals allow users to post their queries and answer the questions. These CQA platforms help experts share their knowledge and new users solve their queries. Initially, these community question-answering platforms focused on providing the most relevant answers to the users’ queries Zhang et al. (2021b)Bachschi et al. (2020). As the CQA platforms keep growing, they not only focus on providing the relevant answers but also on quickly delivering such answers Yazdaninia et al. (2021)Wang et al. (2013). The growth and maturity of such platforms is contingent on a number of variables, including the “quality of questions and answers", the “response time of queries", “user involvement", and “user activity" Moutidis and Williams (2021). The users with expertise in specific domains significantly help the platform’s growth. Based on the users’ activeness on the platform and the importance of their posts, they have been rewarded reputation scores, badges, and additional privileges. In most such CQA platforms driven by a reputation system, the platform’s popularity depends on the response time Bhat et al. (2014) of the question (i.e., time to receive the first answer). So, the users who want to post a query may also want to know the time by which they can expect a response to their question. StackExchange [^1] https://stackexchange.com/ is one of the CQA platforms where the users can ask questions and receive answers from other users. An analysis of such a platform’s dynamics helps the community maintainers adapt and design the framework to address various CQA-related problems Anderson et al. (2012) Mondal et al. (2021) Zhang et al. (2021a) Wang et al. (2018b) Hazra et al. (2023b) Hazra et al. (2023a) Hazra et al. (2021). In this paper, we examine the dynamics of different CQA platforms relating the usual response time with other factors of the portals. We conduct an empirical study on six CQA platforms – Mathematics, Software Engineering, English, Game Development, Chemistry, and Money. To analyze the dynamics of the diverse community question-answering platforms, we look into various factors mentioned here – (1) Characterize the questions in terms of textual information, linguistic characteristics, question tags, votes received by the questions etc. We obtain correlations of these features with the time required to respond to a question. (2) Characterizing the users of the platform in terms of their reputation & activity in the platforms. Further, we build Asker-Answerer graph (AAG) where the nodes are the users and a link is created if the question from a user is answered by another user. Subsequently, we conduct in depth analysis of the characteristics of the common users present across all the platforms. (3)For all the platforms, we use metadata of the questions, question structure and user interaction as features to build a standard machine learning model for classifying the newly posted questions which are having fast answer. The majority of preceding studies Wang et al. (2018a) scrutinize the correlation between diverse question-related factors and response time. The novelty of our current research resides in our endeavor to comprehend how various inherent network-related features (reflected in the asker-answerer graph) coincide with question-related features in relation to response time. In this study, we make the following observations – (1) Short questions get faster responses. Chemistry and Software Engineering questions are harder. More tags mean slower replies. Math, Software Engineering, and English answers come in an hour, others in a day. In Software Engineering and English, many questions get $\geq 3$ answers.(2) Most users ( 60%-70%) on all platforms stay inactive. Domains like Money, Game Development, and Chemistry have more active users than Mathematics, Software Engineering, and English. (3) XGBoost and MLP perform best at classifying new questions on most platforms. The second best is usually the random forest model.